OpenStudy (anonymous):

Write the equation of the line that is parallel to the line 3x − y = −3 and passes through the point (4, −2). y = −one thirdx − 6 y = −one thirdx − 14 y = 3x − 6 y = 3x − 14

OpenStudy (anonymous):

Work out the gradient of the first line. If you put the y onto the other side and bring the -3 across then you should have y=3x+3. The gradient of that is 3 and to have a parrallel line the gradient must be the same so again it's 3x. Then use the formula to calculate the equation of a line. y-y1=m(x-x1) So you have y + 2 = 3(x-4) So y=3x-6

OpenStudy (abb0t):

Find the slope, m \[m = \frac{ y_2 - y_1 }{ x_2 - x_1 }\] then use: \[(y-y_1) = m (x-x_1) \]

OpenStudy (anonymous):

thanks